| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  |  
| 
  22  
 | 
Sat Mar 29 15:51:46 2025 | 
CG, JG, MC, ID | General | DAQ setup | ********************************************
 
** Setup of DAQ x86l-131 (outside pocket) 
**
 | 
   |  | 
  23  
 | 
Sat Mar 29 17:19:46 2025 | 
CG, JG, MC, ID | General | Sample data file | Data files written to /mnt/raw.data/e0203_e0018.
 
 
From MBS computers:
 | 
   |  | 
  24  
 | 
Sat Mar 29 17:53:13 2025 | 
CG, ID, MC | Electronics | Bias remote control | The MHV modules are set up for remote control. 
This can be accessed via the mesycontrol 
software running on x86l-128.
 | 
   |  | 
  25  
 | 
Mon Mar 31 17:55:16 2025 | 
JG | DAQ | DAQ overview | The DAQ is running of 4 different nodes/PCs, 
that can be accessed by ssh using the user 
litv-exp.
 | 
   |  | 
  26  
 | 
Mon Mar 31 18:21:37 2025 | 
JG | DAQ | ESR scaler mapping | This is the list of scaler channels used at 
the VME ESR DAQ on x86l-204:
 
 | 
   |  | 
  27  
 | 
Mon Mar 31 18:35:46 2025 | 
JG | DAQ | lmd files on ESR NAS | The files of each MBS node or the timesorter 
are written to the ESR NAS by the RFIO server.
 
The server is running on the virtual machine | 
   |  | 
  28  
 | 
Tue Apr  1 17:18:10 2025 | 
ID, MC, CJG | Detectors | Both PLEIADES detectors in beam | We started inserting the inside and outside 
detector, see
 
https://elog.gsi.de/esr/E0203_E0018/46
 | 
   |  | 
  30  
 | 
Tue Apr  1 22:41:52 2025 | 
CG | Detectors | Climbing leakage currents | The experiment has been running stably during 
the afternoon.
 
DAQ rates have remained constant and the | 
    |  | 
  31  
 | 
Tue Apr  1 23:00:41 2025 | 
CG | General | Run stopped and detectors moved back | Due to the continued rise in leakage currents, 
I have stopped the run, moved the detectors 
completely back and turned off the 
 | 
   |  | 
  32  
 | 
Thu Apr  3 13:09:28 2025 | 
JG | DAQ | 194Pb MBS settings | The settings of the MBS daq during the first 
data taking with fragments around 194Pb (W.Korten 
experiment) was not entirely correct:
 | 
   |  | 
  33  
 | 
Thu Apr  3 13:24:31 2025 | 
CG | Detectors | Bias problem diagnosing | During an access break to the ESR, I made 
a few tests with the bias scheme to see if 
we could better understand some of the weird | 
   |  | 
  34  
 | 
Thu Apr  3 14:53:13 2025 | 
ID | Detectors | Link to leakage current Google doc | Please enter the leakage currents from the 
mesycontrol every ~1h into this Google document:
 
 | 
   |  | 
  35  
 | 
Thu Apr  3 17:37:04 2025 | 
CJG, GL, MC | General | Detectors in position, taking data | The inner detector was moved into position.
 
The device control for the outer detector 
is not responding and I am not an expert/don't | 
   |  | 
  36  
 | 
Fri Apr  4 16:12:36 2025 | 
CJG, GL, MC, JG, ID | General | Moving PLEIADES in | 186Hf now stored stably in the ring.
 
 
Will move both detectors in to position.
 | 
   |  | 
  37  
 | 
Fri Apr  4 16:27:11 2025 | 
CJG, GL, MC, JG, ID | General | PLEIADES in position | Inner PLEIADES detector moved into position 
at +2mm.
 
Again, the outside pocket cannot be moved.
 | 
   |  | 
  38  
 | 
Fri Apr  4 21:15:13 2025 | 
CJG, GL, MC, ID | General | Changes to setup | After talking to T. Davinson this afternoon 
a few things became clearer.
 
 | 
   |  | 
  39  
 | 
Sat Apr  5 12:28:20 2025 | 
CG, GL, ID | General | Detector moved in | Inside pocket moved in to 2mm
 
 
run0024 started @ 12:27:50 | 
   |  | 
  40  
 | 
Sat Apr  5 12:38:49 2025 | 
CG | General | Current status | Given that I hope we are now moving towards 
stable operation, here is the current status 
of detectors for run0024 on...
 | 
   |  | 
  41  
 | 
Sat Apr  5 12:54:20 2025 | 
CG, MC | Detectors | Bias curves | Bias curves were completed for most detectors 
yesterday evening.
 
 | 
       |  | 
  42  
 | 
Sat Apr  5 13:20:10 2025 | 
ID, CG, GL | General | MHV-4 max current | The leakage currents are 6-7 micro-amps for 
the Si pads at the moment. The max current 
for the MHV-4 is 20 microamps PER CHANNEL, | 
   |   
 |